Consider a restaurant where ovens are a fixed input and workers are variable inputs. Assume labor is the only variable cost for the business. The restaurant has a fixed cost of $50 per day and pays each worker $90 per day.

Fill in the blanks to complete the Marginal Physical Product of Labor column for each worker and the Marginal Cost column at each level of labor. (Hint: Marginal cost is the change in total cost divided by the change in the quantity of output. You can calculate it here by dividing the increase in total cost from hiring one more worker by the marginal physical product from hiring one more worker.)

Labor

Quantity of Output

Marginal Physical Product of Labor

Total Cost

Marginal Cost

(Workers)

(Pizzas per day)

(Pizzas per day)

(Dollars per day)

(Dollars per pizza)

0 0 50
1 10 140
2 25 230
3 45 320
4 75 410
5 100 500
6 115 590

True or False: Marginal physical product and marginal cost move in the same direction.

True

False
Answer #1

Answer:

Labor

Quantity of output

Marginal physical product of Labor = Change in Quantity of output

Total Cost(TC) = 50+ (90)L

Marginal Cost=

(Change in TC/ Change in Quantity of output)

0

0

-

50

-

1

10

10-0 = 10

140

(140 – 50) / 10 = 9

2

25

25 – 10 = 15

230

(230 – 140) / 15 = 6

3

45

45- 25 = 20

320

(320 – 230) / 20 = 4.5

4

75

75 – 45 = 30

410

(410 – 320) / 30 = 3

5

100

100 – 75 = 25

500

(500 – 410) / 25 = 3.6

6

115

115 – 100 = 15

590

(590 – 500)/ 15 = 6

Marginal physical product and marginal cost move in the same direction.

False.

When Marginal physical product of Labor rises the marginal Cost decreases.